OpenStudy (anonymous):

Tickets to a movie cost $9 for each adult and $4.50 for each child. Molly purchased 11 tickets for $72. How many adult and how many child tickets did she purchase? Molly purchased ___ adult tickets and ___ child tickets.

OpenStudy (aripotta):

so this a system of equations. let's make our equations :) i'm gonna let a stand for adult and c for child we know she bought 11 tickets. so a + c = 11 and we also know she spent a total of $72 and adult tickets cost $9 and child tickets cost $4.50. so 9a + 4.5c = 72 so our system is a + c = 11 9a + 4.5c = 72 making sense so far?

OpenStudy (triciaal):

@kewlkatcutie you need to answer before we can move on. This is an interactive site.

OpenStudy (anonymous):

sorry , I was afk

OpenStudy (aripotta):

no worries :)

OpenStudy (anonymous):

yeah, it makes sense

OpenStudy (aripotta):

great. now i'd use the substitution method to solve this. use the first equation to solve for a in terms of c

OpenStudy (aripotta):

then substitute that for a in the second equation

OpenStudy (aripotta):

then solve for c

OpenStudy (anonymous):

Sorry for the late response , but I'm not sure how the substitution method works ?

OpenStudy (aripotta):

ok so the first step of the substitution method is to solve for a variable in terms of the other. so we're going to take the first equation, a + c = 11, and solve for a

OpenStudy (aripotta):

so a = ?

OpenStudy (anonymous):

5 ?

OpenStudy (aripotta):

well, i meant in terms of c. so if a + c = 11, then a = 11 - c

OpenStudy (aripotta):

then we can substitute 11 - c for a in the second equation. 9a + 4.5c = 72 9(11 - c) + 4.5c = 72 now we can actually solve for c

OpenStudy (aripotta):

well i just worked it out all the way, and yes a = 5 :)

OpenStudy (anonymous):

Is it 5 adults and 6 children ?

OpenStudy (aripotta):

yes. it seems as though you didn't even need help tho lol :P